OpenStudy (anonymous):

the numbers of dimes i have is 10 more than twice the number of nickels. the value of my coins is 9$. how many coins do i have

OpenStudy (anonymous):

170 COINS THAT HAVING 160 NICKLES AND TEN DIMES EQUAL TO 9 DOLLARS im not sure though tricky question

OpenStudy (anonymous):

Let's set up a system of two linear equations. The first equation will represent the number of coins we have, and the second will represent the value of the coins. If the number of dimes I have is 10 MORE than twice the number the number of nickels, than I have\[D = 10 + 2N\] I have $9 worth of coins. \[0.1 \cdot D + 0.05 N = 9\]

OpenStudy (anonymous):

thanks you guys! :)